# VidLaDA: Bidirectional Diffusion Large Language Models for Efficient Video Understanding

### v1

@@ B.1 Autoregressive-Based Video Large Language Models
B.2 Diffusion and Non-Autoregressive Language Models
Diffusion models have achieved remarkable success in continuous data generation (Peebles & Xie, 2023; Polyak et al., 2024; Li et al., 2023b). Recently, substantial progress has been made in adapting diffusion to discrete language modeling, including diffusion in continuous embedding space (Li et al., 2022) and masked/discrete formulations such as D3PM (Ho et al., 2020) and SSD-LM (Han et al., 2023). More recent large-scale Diffusion Language Models (DLMs), e.g., LLaDA (Nie et al., 2025), LLaDA-MoE (Zhu et al., 2025), LLaDA2 (Bie et al., 2025), and Dream (Ye et al., 2025c), demonstrate competitive generation quality to strong autoregressive LLMs while enabling parallel decoding via iterative denoising.
In multimodal settings, several works extend DLMs to vision-language tasks by conditioning on spatiotemporal tokens and largely retaining the standard MLLM pipeline (vision encoder $\rightarrow$ projector $\rightarrow$ language backbone), where diffusion is introduced mainly by swapping the AR backbone with a masked-denoising Transformer (e.g., LLaDA-V (You et al., 2025), Dream-VL (Ye et al., 2025a), and concurrent efforts (Li et al., 2025b; Cheng et al., 2025; Xin et al., 2025)). While these studies validate the feasibility of diffusion-based VLMs, they typically provide limited task- and modality-specific motivation or analysis for “Why DLM is more suitable than AR-Based model in image/video understanding?”, and they do not explicitly target the spatiotemporal challenges unique to video understanding.
VidLaDA bridges diffusion modeling and video understanding in a more video-oriented manner. We first analyze and justify why bidirectional DLM is suitable for spatiotemporal understanding in videos, beyond treating diffusion as a drop-in replacement for AR decoding. Building on this motivation, we present (to the best of our knowledge) the first DLM-based Video LLM that demonstrates strong performance on complex video understanding benchmarks. Finally, to make diffusion inference practical under long-form video contexts with massive video tokens, we introduce MARS-Cache, a multimodal, structure-aware caching framework that significantly reduces redundant computation during decoding.
Appendix C Proof
C.1 Proof of Proposition 3.1
Consider a sequence $X=\{x_{1},\dots,x_{T}\}$ processed by a causal attention mechanism. The attention score $A_{i,j}$ (the attention $x_{i}$ pays to $x_{j}$) is valid only if $j\leq i$. This constraint is enforced by a lower-triangular causal mask $M\in\{0,-\infty\}^{T\times T}$.
We define the visibility frequency $\mathcal{C}(x_{j})$ of a token $x_{j}$ as the number of times it acts as a key/value for other tokens in the sequence during a forward pass:
$\mathcal{C}(x_{j})=\sum_{t=1}^{T}\mathbb{I}(j\leq t)=T-j+1,$ (11)
@@ Appendix C Proof
For the initial token $x_{1}$, $\mathcal{C}(x_{1})=T$, meaning it is visible to and attended by every token in the sequence. In contrast, for a later token $x_{k}$ (where $k\gg 1$), $\mathcal{C}(x_{k})$ is significantly smaller.
@@
Because the softmax function $\sigma(x_{j})=e^{x_{j}}/\sum_{k}e^{x_{k}}$ forces weights to sum to 1, the model requires a stable token to absorb excess attention mass when current semantic features are ambiguous (Xiao et al., 2024). Since $x_{1}$ is the only universally accessible token with maximum visibility frequency, the model learns an inductive bias to utilize $x_{1}$ as a specialized asymmetric receptive field.
@@
For the visual sequence, since the index $j=1$ corresponds to a specific spatial location (e.g., top-left corner) due to rasterization, the learned bias induces a spatially non-uniform importance distribution: the visual start token is disproportionately weighted solely due to its position in the input sequence, thereby shaping the early receptive field.
C.2 Proof of Proposition 3.2
We model the information flow within the VLM as a Markov chain with the $M$-layer ViT and the $N$-layer LLM. Let the ViT map the original visual inputs $\mathcal{V}$ to an intermediate sequence of features $H^{(M)}=\{h_{1},\dots,h_{T}\}$. Since the encoder is bidirectional, $H^{(M)}$ captures global visual context. However, the subsequent processing in the $N$-layer LLM introduces a divergence.
Consider the generation of the latent representation $z_{t}$ at step $t\in\{1,\dots,T\}$ in the final layer.
According to the Data Processing Inequality, for any Markov chain $X\to Y\to Z$, we have $I(X;Z)\leq I(X;Y)$.
For bidirectional model $\mathcal{M}_{\text{bidi}}$, the generation of $z_{t}^{\text{bidi}}$ can access the entire intermediate feature set $H^{(M)}$ via full attention. The dependency is $\mathcal{V}\to H^{(M)}\to z_{t}^{\text{bidi}}$. Thus, the information capacity is bounded by the total encoded information:
$I(z_{t}^{\text{bidi}};\mathcal{V})\leq I(H^{(M)};\mathcal{V})$ (12)
For AR model $\mathcal{M}_{\text{uni}}$, due to the unidirectional causal mask, the generation of $z_{t}^{\text{uni}}$ is restricted to the current and past features $H_{\leq t}^{(M)}=\{h_{1},\dots,h_{t}\}$. The dependency is $\mathcal{V}\to H_{\leq t}^{(M)}\to z_{t}^{\text{uni}}$. Thus, the bound is:
$I(z_{t}^{\text{uni}};\mathcal{V})\leq I(H_{\leq t}^{(M)};\mathcal{V})$ (13)
Visual information is spatially distributed. The intermediate features $H^{(M)}$ are lossy compressions of $\mathcal{V}$ with spatial specialization. Consequently, the future tokens $H_{>t}^{(M)}$ contain independent information about $\mathcal{V}$ that is not fully captured by $H_{\leq t}^{(M)}$. This implies a strictly positive conditional mutual information:
$\small\Delta I_{t}=I(H^{(M)};\mathcal{V})-I(H_{\leq t}^{(M)};\mathcal{V})=I(H_{>t}^{(M)};\mathcal{V}\mid H_{\leq t}^{(M)})>0$ (14)
Therefore, for any $t<T$, the latent representation in the AR model has a strictly lower information upper bound than that of the bidirectional model.
We simply assume the final response $R$ is generated based on the probability $P(R\mid Z)$. The quality of $R$ in VLM is constrained by the joint mutual information of the entire sequence $I(R;\mathcal{V})$. Summing the pointwise information capacities (assuming independence for the upper bound estimation):
$\sum_{t=1}^{T}\sup I(z_{t}^{\text{bidi}};\mathcal{V})-\sum_{t=1}^{T}\sup I(z_{t}^{\text{uni}};\mathcal{V})=\sum_{t=1}^{T-1}\Delta I_{t}>0$ (15)
This confirms that the latent sequence $Z^{\text{uni}}$ essentially discards the future visual context during the deep semantic abstraction process (layers $M+1$ to $M+N$), leading to an irreversible loss of visual understanding capability compared to $Z^{\text{bidi}}$.
Appendix D Method Details
D.1 Training Pipeline
D.1.1 Data Composition
@@
To remedy the lack of minute-scale, long-horizon temporal reasoning in existing video understanding training data, we construct a corresponding dataset through a robust construction pipeline. Utilizing FineVideo (Farré et al., 2024) as the seed source, we curate a high-quality dataset specifically emphasizing videos ranging from 2 to 30 minutes. Our pipeline leverages Deepseek-V3.1 (Liu et al., 2024a) as the LLM for instruction synthesis and text-based filtering, and Qwen3-VL-235B-A22B (Bai et al., 2025a) as the VLM for visual consistency verification. The pipeline consists of four distinct stages:
@@
Quality Assurance via Consistency Voting. To guarantee label reliability and minimize hallucinations, we employ a self-consistency voting strategy. The VLM generates responses for each sample three times under varying temperatures. An LLM evaluator then compares these responses against the ground truth. Only samples where the VLM achieves a consistency consensus (i.e., at least 2 out of 3 responses match the ground truth) are retained for the final training set.
Beyond the newly curated long-form video data, we integrate established high-quality benchmarks to ensure robust performance across diverse modalities. Specifically, we adopt LLaVA-Video-178K (Zhang et al., 2024b) as the foundational source for short-to-medium video instruction tuning, ensuring coverage of rich temporal dynamics and general video understanding. Furthermore, to maintain and enhance fine-grained spatial reasoning capabilities, we incorporate the MAmmoTH-VL (Guo et al., 2025a) dataset for image-based instruction following. This holistic data composition, which spans static images, short clips, and our proposed long-form videos, forms the basis of our unified training curriculum.
D.1.2 Multi-Stage Training Strategy
Table 3: Detailed Specifications of the Multi-Stage Training Curriculum. We outline the data composition, temporal resolution scaling, and optimization configurations across three stages, transitioning from short-clip alignment to long-form video reasoning.
@@ D.1.2 Multi-Stage Training Strategy
Stage 1: Short-Clip Temporal Pre-Training. The primary objective of the initial stage is to equip the static VLM with fundamental temporal perception capabilities. We utilize a large-scale collection of 1.8M short videos and images, with durations spanning from approximately 10 seconds to 3 minutes, sourced from diverse datasets. In this phase, the model is trained with a temporal resolution of 32 frames and a maximum sequence length of 8K tokens. We perform full-parameter fine-tuning on the Vision Transformer (ViT), Projector (MLP), and LLM backbone to align the vision encoder’s temporal features with the language space using learning rates of 2e-6, 1e-5, and 1e-5, respectively.
@@
D.2 Chain of Thought Prompting Details
@@
The specific prompt templates used for each stage are detailed in Tables 10 through 13.
Appendix E More Experimental Results
E.1 The Role of Video Dataset Duration
This section complements Section D.1.1 by analyzing how the duration distribution of training videos affects long-context video understanding. We follow the baseline training recipe in Table 3 and construct two subsets by duration: 1min-3min and 2min-30min. For each range, we uniformly sample 20K instances and train with the same hyperparameters and input configuration, so the only variable is the duration distribution.
Table 4 shows that both duration-restricted subsets underperform the baseline on LongVideoBench and LVBench, suggesting that reducing diversity by sampling from a single duration range can hurt generalization. Meanwhile, the 2min-30min subset achieves clear gains over both the baseline and the 1min-3min subset on long-video robustness, most notably on Video-MME (62.1 to 64.2) and also improves MVBench (58.2 to 58.5). In contrast, the 1min-3min subset consistently degrades across long-video and perception benchmarks.
Long-duration data (2min-30min) is beneficial for long-context robustness (especially Video-MME), likely because it exposes the model to longer-range temporal dependencies. However, using only long-duration videos can slightly reduce performance on some long-video QA benchmarks (e.g., LongVideoBench 62.0 to 61.6), indicating that duration-only filtering may trade off data diversity and task coverage. Overall, the results support using long-duration data as an important component, while keeping mixed-duration data.
@@ E.1 The Role of Video Dataset Duration
2min-30min 64 61.6 43.5 58.5 64.2
Table 4: Comparison with Different Video Dataset Duration.
E.2 Multi-Stage Training
This section empirically validates the efficacy of the multi-stage curriculum detailed in Section D.1.2 by evaluating checkpoints at the end of each training stage. Table 5 presents the quantitative progression of the model performance.
@@ E.2 Multi-Stage Training
3 64 46.6 61.4 44.7 64.5 59.4 69.2 53.4 64.2 57.9
Table 5: Stage Performance Comparison.
E.3 CoT Inference Visualizations

### v2

@@ C.1 Autoregressive-Based Video Large Language Models
C.2 Diffusion and Non-Autoregressive Language Models
Diffusion models have achieved remarkable success in continuous data generation (Peebles & Xie, 2023; Polyak et al., 2024; Li et al., 2023b; Wang et al., 2024b; Gan et al., 2025). Recently, substantial progress has been made in adapting diffusion to discrete language modeling, including diffusion in continuous embedding space (Li et al., 2022) and masked/discrete formulations such as D3PM (Ho et al., 2020) and SSD-LM (Han et al., 2023). More recent large-scale Diffusion Language Models (DLMs), e.g., LLaDA (Nie et al., 2025), LLaDA-MoE (Zhu et al., 2025), LLaDA2 (Bie et al., 2025), and Dream (Ye et al., 2025c), demonstrate competitive generation quality to strong autoregressive LLMs while enabling parallel decoding via iterative denoising. In multimodal settings, several works extend DLMs to vision-language tasks by conditioning on spatiotemporal tokens and largely retaining the standard MLLM pipeline (vision encoder $\rightarrow$ projector $\rightarrow$ language backbone), where diffusion is introduced mainly by swapping the AR backbone with a masked-denoising Transformer (e.g., LLaDA-V (You et al., 2025), Dream-VL (Ye et al., 2025a), and concurrent efforts (Li et al., 2025b; Cheng et al., 2025; Xin et al., 2025)).
While these studies validate the feasibility of DLM-based MLLMs and acknowledge the benefits of bidirectional context, our work differentiates itself by offering a formalized theoretical analysis (Propositions 3.1 and 3.2) and rigorous empirical analysis. We explicitly target the spatiotemporal challenges unique to video understanding to demonstrate “Why DLM is structurally more suitable than AR in image/video understanding?”, beyond general generative capabilities. VidLaDA bridges diffusion modeling and video understanding in a more video-oriented manner. We first analyze and justify why bidirectional DLM is suitable for spatiotemporal understanding in videos, beyond treating diffusion as a drop-in replacement for AR decoding. Building on this motivation, we present (to the best of our knowledge) the first DLM-based Video LLM that demonstrates strong performance on complex video understanding benchmarks. Finally, to make diffusion inference practical under long-form video contexts with massive video tokens, we introduce MARS-Cache, a multimodal, structure-aware caching framework that significantly reduces redundant computation during decoding.
Appendix D Proof
D.1 Proof of Proposition 3.1
Consider a sequence ${\bm{X}}=\{{\bm{x}}_{1},\dots,{\bm{x}}_{|{\bm{X}}|}\}$ processed by a causal attention mechanism. The attention score ${\bm{A}}_{i,j}$ (the attention ${\bm{x}}_{i}$ pays to ${\bm{x}}_{j}$) is valid only if $j\leq i$. This constraint is enforced by a lower-triangular causal mask ${\bm{M}}\in\{0,-\infty\}^{T\times T}$.
We define the visibility frequency $\mathcal{C}({\bm{x}}_{j})$ of a token ${\bm{x}}_{j}$ as the number of times it acts as a key/value for other tokens in the sequence during a forward pass:
$\mathcal{C}({\bm{x}}_{j})=\sum_{t=1}^{|{\bm{X}}|}\mathbb{I}(j\leq t)=|{\bm{X}}|-j+1,$ (11)
@@ Appendix D Proof
For the initial token ${\bm{x}}_{1}$, $\mathcal{C}({\bm{x}}_{1})=|{\bm{X}}|$, meaning it is visible to and attended by every token in the sequence. In contrast, for a later token ${\bm{x}}_{k}$ (where $k\gg 1$), $\mathcal{C}({\bm{x}}_{k})$ is significantly smaller.
@@
Because the softmax function $\sigma({\bm{x}}_{j})=e^{{\bm{x}}_{j}}/\sum_{k}e^{{\bm{x}}_{k}}$ forces weights to sum to 1, the model requires a stable token to absorb excess attention mass when current semantic features are ambiguous (Xiao et al., 2024). Since ${\bm{x}}_{1}$ is the only universally accessible token with maximum visibility frequency, the model learns an inductive bias to utilize ${\bm{x}}_{1}$ as a specialized asymmetric receptive field.
@@
For the visual sequence ${\bm{E}}^{\mathcal{V}}$}, since the index $j=1$ corresponds to a specific spatial location (e.g., top-left corner) due to rasterization, resulting in $\mathcal{C}({\bm{e}}^{\mathcal{V}}_{1})=|{{\bm{E}}^{\mathcal{V}}}|$, the learned bias induces a spatially non-uniform importance distribution: the visual start token is disproportionately weighted solely due to its position in the input sequence, thereby shaping the early receptive field.
D.2 Proof of Proposition 3.2
This section establishes the formal proof regarding the information capacity gap between the Autoregressive (AR) and Bidirectional (Bidi) decoding paradigms.
Let $\mathcal{V}$} be the random variable representing the input video. We denote the encoded visual feature sequence produced by the Vision Transformer (ViT) as ${{\bm{E}}^{\mathcal{V}}}=\{{\bm{e}}^{\mathcal{V}}_{1},\dots,{\bm{e}}^{\mathcal{V}}_{|{{\bm{E}}^{\mathcal{V}}}|}\}$, assuming it preserves the global information of $\mathcal{V}$}. We model the decoding process as a Markov chain where ${\bm{z}}_{t}$ is the latent state at the final layer of the LLM at position $t$. The Bidirectional paradigm utilizes full self-attention, characterized by the Markov chain ${\mathcal{V}}\to{{\bm{E}}^{\mathcal{V}}}\to{\bm{z}}_{t}^{\text{bidi}}$. Conversely, the Autoregressive paradigm (AR) is conditioned strictly on the history ${{\bm{E}}^{\mathcal{V}}}_{\leq t}=\{{\bm{e}}^{\mathcal{V}}_{1},\dots,{\bm{e}}^{\mathcal{V}}_{t}\}$, characterized by ${\mathcal{V}}\to{{\bm{E}}^{\mathcal{V}}}_{\leq t}\to{\bm{z}}_{t}^{\text{AR}}$.
To quantify the disparity, we define the cumulative pointwise information capacity as $\mathcal{K}\triangleq\sum_{t=1}^{T}I({\bm{z}}_{t};{\mathcal{V}})$. This metric represents the aggregate source information accessible across the generated sequence. Our analysis rests on the following two theoretical assumptions.
Assumption D.1.
Visual information is spatially distributed and the encoder is non-redundant. Specifically, for any $t<T$, the future visual features ${{\bm{E}}^{\mathcal{V}}}_{>t}$ contain non-zero conditional mutual information about $\mathcal{V}$}:
$I({{\bm{E}}^{\mathcal{V}}}_{>t};\mathcal{V}\mid{{\bm{E}}^{\mathcal{V}}}_{\leq t})>0.$
Assumption D.2.
The LLM decoder possesses sufficient expressivity such that the information bottleneck is dominated by architectural visibility constraints. Consequently, the supremum of the mutual information is bounded by the Data Processing Inequality (DPI):
$\sup I({\bm{z}}_{t}^{\text{bidi}};{\mathcal{V}})=I({{\bm{E}}^{\mathcal{V}}};{\mathcal{V}}),$ (12)
$\sup I({\bm{z}}_{t}^{\text{AR}};{\mathcal{V}})=I({{\bm{E}}^{\mathcal{V}}}_{\leq t};{\mathcal{V}}).$ (13)
Proceeding with the derivation, we compare the summation terms for the two paradigms. Applying the DPI to the Markov chains defined above and invoking Assumption D.2, the information capacity gap at step $t$ is given by:
$\Delta_{t}\triangleq\sup I({\bm{z}}_{t}^{\text{bidi}};\mathcal{V})-\sup I({\bm{z}}_{t}^{\text{AR}};\mathcal{V})=I({{\bm{E}}^{\mathcal{V}}};\mathcal{V})-I({{\bm{E}}^{\mathcal{V}}}_{\leq t};\mathcal{V}).$ (14)
Using the chain rule for Mutual Information, we decompose the joint information of sequence ${\bm{E}}^{\mathcal{V}}$} as:
$I({{\bm{E}}^{\mathcal{V}}};\mathcal{V})=I({{\bm{E}}^{\mathcal{V}}}_{\leq t};\mathcal{V})+I({{\bm{E}}^{\mathcal{V}}}_{>t};\mathcal{V}\mid{{\bm{E}}^{\mathcal{V}}}_{\leq t}).$ (15)
Substituting this yields:
$\Delta_{t}=I({{\bm{E}}^{\mathcal{V}}}_{>t};\mathcal{V}\mid{{\bm{E}}^{\mathcal{V}}}_{\leq t}).$ (16)
Under Assumption D.1, it holds that $\Delta_{t}>0$ for all $t<T$. Finally, summing over the sequence length $T$ reveals a strict gap in cumulative capacity:
$\sum_{t=1}^{T}\sup I({\bm{z}}_{t}^{\text{bidi}};\mathcal{V})-\sum_{t=1}^{T}\sup I({\bm{z}}_{t}^{\text{AR}};\mathcal{V})$ $=\sum_{t=1}^{T-1}\Delta_{t}+\Delta_{T}$
$=\sum_{t=1}^{T-1}I({{\bm{E}}^{\mathcal{V}}}_{>t};\mathcal{V}\mid{{\bm{E}}^{\mathcal{V}}}_{\leq t})+0$
$>0.$
Appendix E Method Details
E.1 Training Pipeline
E.1.1 Data Composition
@@
To remedy the lack of minute-scale, long-horizon temporal understanding in existing video instruction-tuning data, we construct a corresponding dataset through a robust construction pipeline. Utilizing FineVideo (Farré et al., 2024) as the seed source, we curate a high-quality dataset specifically emphasizing videos ranging from 2 to 30 minutes. Our pipeline leverages Deepseek-V3.1 (Liu et al., 2024a) as the LLM for instruction synthesis and text-based filtering, and Qwen3-VL-235B-A22B (Bai et al., 2025a) as the MLLM for visual consistency verification. The pipeline consists of four distinct stages:
@@
Quality Assurance via Consistency Voting. To guarantee label reliability and minimize hallucinations, we employ a self-consistency voting strategy. The MLLM generates responses for each sample three times under varying temperatures. An LLM evaluator then compares these responses against the reference answers derived from the instruction synthesis phase. Only samples where the MLLM achieves a consistency consensus (i.e., at least 2 out of 3 responses match the ground truth) are retained for the final training set.
Beyond the newly curated video data, we integrate established high-quality benchmarks to ensure robust performance across diverse modalities. Specifically, we adopt LLaVA-Video-178K (Zhang et al., 2024b) as the foundational source for short-to-medium video instruction tuning, ensuring coverage of rich temporal dynamics and general video understanding. Furthermore, to maintain and enhance fine-grained spatial reasoning capabilities, we incorporate the MAmmoTH-VL (Guo et al., 2025a) dataset for image-based instruction following. This holistic data composition, which spans static images, short clips, and our proposed parts, forms the basis of our unified training curriculum.
E.1.2 Multi-Stage Training Strategy
Table 4: Detailed Specifications of the Multi-Stage Training Curriculum. We outline the data composition, temporal resolution scaling, and optimization configurations across three stages, transitioning from short-clip alignment to long-form video understanding.
@@ E.1.2 Multi-Stage Training Strategy
Stage 1: Short-Clip Temporal Pre-Training. The primary objective of the initial stage is to equip the static MLLM with fundamental temporal perception capabilities. We utilize a large-scale collection of 1.8M short videos and images, with durations spanning from approximately 10 seconds to 3 minutes, sourced from diverse datasets. In this phase, the model is trained with a temporal resolution of 32 frames and a maximum sequence length of 8K tokens. We perform full-parameter fine-tuning on the Vision Transformer (ViT), Projector (MLP), and LLM backbone to align the vision encoder’s temporal features with the language space using learning rates of 2e-6, 1e-5, and 1e-5, respectively.
@@
E.2 Chain of Thought Prompts Details
@@
The specific prompt templates used for each stage are detailed in Tables 12 through 15.
Appendix F More Experimental Results
F.1 The Role of Video Dataset Duration
This section complements Appendix E.1.1 by analyzing how the duration distribution of training videos affects video understanding. We follow the training recipe in Table 11, adopt the learning rate and trainable modules from Stage 3 of Table 4, and use the Stage 2 checkpoint as the baseline. We then construct two subsets by duration: 1min-3min and 2min-30min. For each range, we uniformly sample 20K samples and train with the same hyperparameters and input configuration, so the only variable is the duration distribution.
As shown in Table 5, the 1min-3min subset consistently underperforms the baseline, particularly on LongVideoBench (62.0 to 59.8). This indicates that restricting training to short clips limits the temporal receptive field, causing the model to overfit to short-term dependencies and degrade on long-range reasoning tasks. Conversely, the 2min-30min subset excels in temporal modeling, surpassing the baseline on Video-MME (62.1 to 64.1), LVBench, and MVBench. The improvement on Video-MME confirms that exposure to extended temporal spans is essential for learning to bridge distant visual cues, a capability not achievable through short clips alone.
However, the 2min-30min subset slightly trails the baseline on LongVideoBench (62.0 to 61.6). This may suggest a trade-off: long-duration data enhances temporal depth, while short-duration data contributes to local semantic perception. Thus, prioritizing long-duration data while retaining a mixed distribution offers the optimal balance for broad generalization.
Table 5: Comparison with Different Video Dataset Duration.
@@ F.1 The Role of Video Dataset Duration
2min-30min 64 61.6 43.5 58.5 64.1
F.2 Multi-Stage Training
This section empirically validates the efficacy of the multi-stage curriculum detailed in Section E.1.2 by evaluating checkpoints at the end of each training stage. Table 6 presents the quantitative progression of the model performance.
@@ F.2 Multi-Stage Training
3 64 46.6 61.4 44.7 64.5 59.4 69.2 53.4 64.1 57.9
F.3 CoT Inference Visualizations